For Taylor Corporation, the working capital at the end of the current year is $10,000 more than the working capital at the end of the preceding year, reported as follows:

Year 2

Year 1

Current assets:

Cash, marketable securities, and receivables

$ 80,000

$ 84,000

Inventories

120,000

    66,000

Total current assets

$200,000

$150,000

Current liabilities

100,000

    60,000

Working capital

$100,000

$ 90,000

Has the current position of Taylor Corporation improved? Explain
Answer #1

current ratio in Year 1 = 150/60 = 2.5

current ratio in Year 2 = 200/100 = 2

so the current ratio has decline indicating weakening in liquidity